OpenStudy (anonymous):

can some one help with this on 7y*8y

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

$7y\cdot\:8y=56y^2$7y· 8y=56y2 Steps $7y\cdot\:8y$7y· 8y $\mathrm{Multiply\:the\:numbers:}\:7\cdot\:8=56$Multiply the numbers: 7· 8=56 $=56yy$=56yy $\mathrm{Apply\:exponent\:rule}:\quad\:a^b\cdot\:a^c=a^{b+c}$Apply exponent rule: ab· ac=ab+c $yy=\:y^{1+1}=\:y^2$yy= y1+1= y2 $=56y^2$=56y​2

OpenStudy (anonymous):

its -7y to the 7th power * 8y to the 4th power

OpenStudy (anonymous):

u did not put that

OpenStudy (phi):

you could use the equation editor (see the button in the lower left area below the input box) \[ \left( -7y \right)^7 \cdot \left(8y\right)^4 \]

OpenStudy (anonymous):

-7y^7(8y^4)

OpenStudy (anonymous):

sorry I got the problem mixed upp

OpenStudy (anonymous):

both are like terms with

OpenStudy (anonymous):

$-7y^7\left(8y^4\right)$−7y7(8y4) $\mathrm{Remove\:parentheses}:\quad\left(a\right)=a$Remove parentheses: (a)=a $=-7y^7\cdot\:8y^4$=−7y7· 8y4 $\mathrm{Multiply\:the\:numbers:}\:7\cdot\:8=56$Multiply the numbers: 7· 8=56 $=-56y^7y^4$=−56y7y4 $\mathrm{Apply\:exponent\:rule}:\quad\:a^b\cdot\:a^c=a^{b+c}$Apply exponent rule: ab· ac=ab+c $y^7y^4=\:y^{7+4}=\:y^{11}$y7y4= y7+4= y11 $=-56y^{11}$=−56y11

OpenStudy (anonymous):

=−56y^11

OpenStudy (anonymous):

thank you